integrate the following integral
∫ sin x sqrt(1 + cos 2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sinx\sqrt{1+cos2x}dx
I = \int sinx\sqrt{2cos^{2}x}dx
I = \int \sqrt{2}sinxcosxdx
I = \frac{1}{\sqrt{2}}\int sin2xdx
I = \frac{-1}{2\sqrt{2}} cos2x+c